# Soft-Label Caching and Sharpening for Communication-Efficient Federated Distillation

### v1

@@ Abstract
Federated Learning (FL) enables collaborative model training across decentralized clients, enhancing privacy by keeping data local. Yet conventional FL, relying on frequent parameter-sharing, suffers from high communication overhead and limited model heterogeneity. Distillation-based FL approaches address these issues by sharing predictions (soft-labels) instead, but they often involve redundant transmissions across communication rounds, reducing efficiency. We propose SCARLET, a novel framework integrating synchronized soft-label caching and an enhanced Entropy Reduction Aggregation (Enhanced ERA) mechanism. SCARLET minimizes redundant communication by reusing cached soft-labels, achieving up to 50% reduction in communication costs compared to existing methods while maintaining accuracy. Enhanced ERA can be tuned to adapt to non-IID data variations, ensuring robust aggregation and performance in diverse client scenarios. Experimental evaluations demonstrate that SCARLET consistently outperforms state-of-the-art distillation-based FL methods in terms of accuracy and communication efficiency. The implementation of SCARLET is publicly available at https://github.com/kitsuyaazuma/SCARLET.
Index Terms:
Federated learning, knowledge distillation, non-IID data, communication efficiency.
@@ I Introduction
Conventional FL methods, such as Federated Averaging (FedAvg) [1], face several fundamental challenges that hinder their scalability and practicality in real-world applications. These challenges include high communication costs from frequent transmission of model parameters, heterogeneous device capabilities and network conditions, convergence difficulties due to non-IID data, and privacy concerns related to the exposure of sensitive information [2]. Addressing these issues is crucial for enhancing the efficiency and robustness of FL systems.
Figure 1: Conceptual overview of distillation-based FL approaches. (a) Traditional distillation-based FL transmits all soft-labels in each communication round. (b) SCARLET requests soft-labels only for data not present in the server cache and synchronizes cache updates between the server and clients, significantly reducing communication costs.
To mitigate these limitations, recent studies have explored knowledge distillation-based FL approaches, which offer several advantages over conventional parameter-sharing methods. Instead of transmitting full model parameters, these approaches share model predictions (soft-labels), effectively reducing communication overhead [3], [4], [5]. They also provide greater flexibility in model architecture, addressing the systems heterogeneity challenge by allowing clients to operate with heterogeneous models of varying computational capacities [6], [7]. Moreover, since knowledge distillation involves sharing only aggregated or processed outputs rather than raw model updates, it enhances privacy protection by minimizing the risk of exposing sensitive training data through shared gradients [8], [9].
However, existing distillation-based FL methods exhibit significant inefficiencies. Specifically, these methods redundantly transmit predictions for identical public dataset samples across multiple communication rounds, despite minimal changes in these predictions. This redundancy results in unnecessary communication overhead. Furthermore, current entropy-based aggregation methods used in distillation-based FL (e.g., ERA in [3]) often lack robustness against variations in data heterogeneity across clients, making parameter tuning challenging and degrading performance in diverse non-IID scenarios.
@@
Our experiments demonstrate that SCARLET consistently outperforms other distillation-based FL methods, achieving target accuracy with reduced communication and faster convergence. This contribution establishes SCARLET as a promising approach for efficient and scalable Federated Learning across heterogeneous client environments.
@@
  • Development of Enhanced ERA: We propose Enhanced ERA, an improvement over traditional ERA, which prevents overfitting in the global model even in weak non-IID settings by dynamically adjusting aggregation sharpness through a controllable hyperparameter.
@@ II-A Federated Learning (FL)
Federated Learning (FL) is a framework for training machine learning models across decentralized clients while preserving data privacy. Since its introduction by [1] with the Federated Averaging (FedAvg) algorithm, FL research has focused on challenges like communication efficiency, non-IID data distributions, and scalability.
Handling heterogeneous data, or non-IID data, where client distributions differ significantly, is a key challenge in FL. Traditional FL methods such as FedAvg are not optimized for these non-IID settings, resulting in slower convergence and reduced accuracy. Various solutions have been proposed, including personalized FL methods like pFedMe [10], which allows client-specific adaptation, and regularization-based approaches such as FedProx [11], which adds a penalty term to local updates to mitigate client drift. SCAFFOLD [12] addresses local model divergence using control variates, and FedNTD [13] employs selective distillation to retain knowledge from “not-true” classes, maintaining a global perspective without extra communication.
@@ II-B Communication-Efficient FL
Quantization reduces the bit precision of model updates, thus lowering data size. FedPAQ [14] applies quantization to compressed client updates for efficiency. FedCOMGATE [15] combines quantization with gradient tracking, aligning local and global updates to support non-IID data while reducing communication.
Sparsification transmits only a subset of gradient elements, focusing on key information. Adaptive Gradient Sparsification [16] dynamically adjusts sparsity based on client needs, using a bidirectional top-k approach that retains key gradients to balance communication efficiency and model accuracy.
Knowledge distillation, which shares distilled knowledge (such as output logits) instead of full parameters, allows model heterogeneity by transmitting only essential information. This approach accommodates model heterogeneity among clients by transferring essential output information rather than the model structure itself. Within distillation methods, approaches like FedDF [17] aggregate client outputs on the server, allowing clients to download an updated model without the need to exchange full parameters. Similarly, FedKD [18] reduces communication costs by employing mutual distillation between a global model and a local model while dynamically compressing exchanged gradients, achieving significant communication savings without sacrificing performance. However, our research focuses on a different class of distillation methods that exclusively exchange model outputs bidirectionally between clients and the server, further enhancing communication efficiency.
@@ II-C Distillation-Based FL
Distillation-based FL reduces the communication overhead from parameter exchange by transmitting only model outputs (logits or soft-labels). Unlike traditional knowledge distillation [19], which compresses models, distillation-based FL enables collaborative learning among heterogeneous clients, managing model heterogeneity while reducing communication.
Some studies use public data in distillation. FedMD [20] enables clients with different models to collaborate by sharing logits derived from a labeled public dataset, allowing clients to refine models without sharing raw data. FD [21] reduces communication by aggregating label-wise average logits normalized by a softmax function from clients’ private data, while DS-FL [3] improves on FD by using unlabeled public data and introducing Entropy Reduction Aggregation (ERA) to handle high-entropy aggregated soft-labels in non-IID settings.
Recent methods further optimize communication, personalization, and non-IID handling. CFD [4] employs soft-label quantization and delta coding to reduce data transmission, and COMET [22] clusters clients with similar data, sharing only soft-labels for improved personalization. Selective-FD [23] filters ambiguous knowledge through selective sharing, using client-side selectors and a server-side selector to handle out-of-distribution samples and high-entropy predictions.
However, existing distillation-based FL methods suffer from redundant communication, as they do not evaluate how significantly soft-labels are updated in each communication round and whether transmitting these updated labels is truly necessary. This limitation motivates our proposed method, SCARLET, which explicitly addresses such inefficiencies by introducing a soft-label caching mechanism to minimize unnecessary transmissions.
@@ II-D FL with Caching Mechanisms
Several studies have proposed FL methods incorporating caching mechanisms. CacheFL [24] employs a caching mechanism that locally stores global models to reduce training time, enabling resource-limited clients to quickly access cached models. FedCache [25] utilizes sample-specific hash mappings and server-side caching to efficiently deliver personalized knowledge without relying on public datasets. FedCache 2.0 [26] replaces logit exchanges with dataset distillation, treating the distilled datasets as a centralized knowledge cache on the server, from which clients download information aligned with their data.
@@ III-A Background and Notations
In FedAvg [1], a standard FL method, each client $k\in[K]$ first updates its model $\mathbf{w}_{k}$ with its private dataset $\mathcal{B}_{k}$ using stochastic gradient descent (SGD). Then, each client uploads its updated model parameters or gradients to the server. The server aggregates these uploaded parameters to obtain the updated global model $\mathbf{w}_{g}$ as follows:
@@
[…]
In DS-FL [3], a distillation-based FL method that serves as the baseline for SCARLET, each client and the server exchange soft-labels instead of model parameters. We assume that each client $k\in[K]$ has not only its labeled private dataset $\mathcal{B}_{k}$ but also the public unlabeled dataset $\mathcal{P}$. After each client updates its model $\mathbf{w}_{k}$ as in FedAvg, it generates local soft-labels $z_{k}^{t}=\sigma(\mathbf{x};\mathbf{w}_{k})$ for $\mathbf{x}\in\mathcal{P}^{t}$, where $\mathcal{P}^{t}$ is unlabeled public dataset used in communication round $t$. The local soft-labels $z_{k}^{t}$ are uploaded from each client to the server, and the server aggregates them to create the global soft-labels $\hat{z}^{t}$ with ERA as follows:
@@
where $\eta_{\mathrm{dist}}$ is the learning rate and $\phi_{\mathrm{dist}}(\cdot,\mathbf{w}_{k}^{t},\cdot)$ is the loss function in the distillation procedure. The loss function $\phi_{\mathrm{dist}}$ is exemplified by the Kullback–Leibler divergence. Along with the clients’ local models, the global model $\mathbf{w}_{g}^{t}$ on the server is also updated with the global soft-labels $\hat{z}_{t}$ and public unlabeled dataset $\mathcal{P}^{t}$.
To avoid ambiguity, we clarify that in this paper, we refer to the outputs shared in distillation-based FL methods as soft-labels rather than logits. Although the term logits is often used in some prior works, such as [3], to refer to the output of the softmax function, it more commonly denotes the unnormalized model output before softmax function. Consequently, we will consistently use the term soft-labels to describe the normalized outputs used in the distillation process, aligning with standard terminology in classification tasks.
@@
$N$ Number of classes for the classification task
@@
Algorithm 1 SCARLET
Server Initialization:
Initialize global model $\mathbf{w}_{g}$ and all client models $\{\mathbf{w}_{k}\}_{k=1}^{K}$
Initialize global cache $\mathcal{C}_{g}$ and all local caches $\{\mathcal{C}_{k}\}_{k=1}^{K}$
Set cache duration $D$
Distribute public dataset $\mathcal{P}$ to all clients $k\in[K]$
for $t=1,\cdots,T$ communication rounds do:
     Server:
     Randomly select subset $\mathcal{P}^{t}\subset\mathcal{P}$
     Compute $\mathcal{I}_{\text{req}}^{t}\leftarrow\{i\in\mathcal{I}^{t}\mid i\notin\text{dom}(\mathcal{C}_{g})\}$
     Send $\mathcal{I}_{\text{req}}^{t}$ to all clients $k\in[K]$
     if $t>1$ then
         Send $\gamma^{t-1},\hat{z}_{\text{req}}^{t-1},\mathcal{I}^{t-1}$ to all clients $k\in[K]$
     end if
     Client $k\in[K]$ (in parallel):
     $\mathbf{w}_{k}^{(t,0)}\leftarrow\mathbf{w}_{k}^{(t-1,E)}$
     if $t>1$ then
         $\hat{z}^{t-1}\leftarrow$ UpdateLocalCache(
                                          $\mathcal{C}_{k}$, $\gamma^{t-1}$, $\hat{z}_{\text{req}}^{t-1}$, $\mathcal{I}^{t-1}$)
         $\mathbf{\tilde{w}}_{k}^{(t,0)}\leftarrow\mathbf{w}_{k}^{(t,0)}$
         for $i=1,\cdots,E_{\text{dist}}$ do:
              $\mathbf{\tilde{w}}_{k}^{(t,i)}\leftarrow\newline \hskip 60.00009pt\mathbf{\tilde{w}}_{k}^{(t,i-1)}-\eta_{\text{dist}}\nabla\phi_{\text{dist}}(\mathcal{P}^{t-1},\mathbf{\tilde{w}}_{k}^{(t,i-1)},\hat{z}^{t-1})$
         end for
         $\mathbf{w}_{k}^{(t,0)}\leftarrow\mathbf{\tilde{w}}_{k}^{(t,E_{d})}$
     end if
     for $j=1,\cdots,E$ do:
         $\mathbf{w}_{k}^{(t,j)}\leftarrow\mathbf{w}_{k}^{(t,j-1)}-\eta\nabla\phi(\mathcal{B}_{k},\mathbf{w}_{k}^{(t,j-1)})$
     end for
     $\mathcal{P}_{k}^{t}\leftarrow\{\mathcal{P}[i]\mid i\in\mathcal{I}_{\text{req}}^{t}\}$
     Send soft-labels $z_{k}^{t}\leftarrow\sigma(\mathcal{P}_{k}^{t},\mathbf{w}_{k}^{(t,E)})$ to server
     Server:
     Set $\mathbf{w}_{g}^{(t,0)}\leftarrow\mathbf{w}_{g}^{(t-1,E)}$
     Aggregate $\{z_{k}^{t}\}_{k\in[K]}$ with 4 to get $\hat{z}_{\text{req}}^{t}$
     $\hat{z}_{\text{cached}}^{t}\leftarrow\{\mathcal{C}_{g}(i)\mid i\in\mathcal{I}^{t}\setminus\mathcal{I}_{\text{req}}^{t}\}$
     $\hat{z}^{t}\leftarrow\hat{z}_{\text{req}}^{t}\cup\hat{z}_{\text{cached}}^{t}$
     for $i=1,\cdots,E$ do:
         $\mathbf{w}_{g}^{(t,i)}\leftarrow\newline \hskip 80.00012pt\mathbf{w}_{g}^{(t,i-1)}-\eta_{d}\nabla\phi_{d}(\mathcal{P}^{t},\mathbf{w}_{g}^{(t,i-1)},\hat{z}^{t})$
     end for
     $\gamma^{t}\leftarrow$ UpdateGlobalCache($\mathcal{C}_{g}$, $\hat{z}^{t}$, $\mathcal{I}^{t}$, $t$, $D$)
end for

### v3

@@ Abstract
Federated Learning (FL) enables collaborative model training across decentralized clients, enhancing privacy by keeping data local. Yet conventional FL, relying on frequent parameter-sharing, suffers from high communication overhead and limited model heterogeneity. Distillation-based FL approaches address these issues by sharing predictions (soft-labels, i.e., normalized probability distributions) instead, but they often involve redundant transmissions across communication rounds, reducing efficiency. We propose SCARLET, a novel framework integrating synchronized soft-label caching and an enhanced Entropy Reduction Aggregation (Enhanced ERA) mechanism. SCARLET minimizes redundant communication by reusing cached soft-labels, achieving up to 50% reduction in communication costs compared to existing methods while maintaining competitive accuracy. Enhanced ERA resolves the fundamental instability of conventional temperature-based aggregation, ensuring robust control and high performance in diverse client scenarios. Experimental evaluations demonstrate that SCARLET consistently outperforms state-of-the-art distillation-based FL methods in terms of accuracy and communication efficiency. The implementation of SCARLET is publicly available at https://github.com/kitsuyaazuma/SCARLET.
@@ I Introduction
Conventional FL methods, such as Federated Averaging (FedAvg) [fedavg], face several fundamental challenges that hinder their scalability and practicality in real-world applications. These challenges include high communication costs from frequent transmission of model parameters, heterogeneous device capabilities and network conditions, convergence difficulties due to non-IID data, and privacy concerns related to the exposure of sensitive information [challenges]. Addressing these issues is crucial for enhancing the efficiency and robustness of FL systems.
(a) Distillation-based FL
(b) SCARLET
Figure 1: Conceptual overview of distillation-based FL approaches. Conventional distillation-based FL (1a) transmits all soft-labels in each communication round. SCARLET (1b) requests soft-labels only for data not present in the server cache and synchronizes cache updates between the server and clients, significantly reducing communication costs.
To mitigate these limitations, recent studies have explored knowledge distillation-based FL approaches, which offer several advantages over conventional parameter-sharing methods. Instead of transmitting full model parameters, these approaches share model predictions (soft-labels [^1] In this paper, we refer to the normalized outputs of the softmax function as soft-labels, and the unnormalized, pre-activation model outputs as logits. A more detailed discussion of this notation is provided in Section III-A.), effectively reducing communication overhead [dsfl], [cfd], [fedgen]. They also provide greater flexibility in model architecture, addressing the systems heterogeneity challenge by allowing clients to operate with heterogeneous models of varying computational capacities [ktpfl], [hbs]. Moreover, since knowledge distillation involves sharing only aggregated or processed outputs rather than raw model updates, it enhances privacy protection by minimizing the risk of exposing sensitive training data through shared gradients [fedad], [fedkdprivacy].
However, existing distillation-based FL methods exhibit significant inefficiencies. Specifically, these methods redundantly transmit predictions for identical public dataset samples across multiple communication rounds, despite minimal changes in these predictions. This redundancy results in unnecessary communication overhead. Furthermore, conventional entropy-based aggregation methods (e.g., ERA in [dsfl]) suffer from a fundamental instability, as their temperature parameter is highly sensitive to input entropy variations. This instability makes robust tuning impractical and degrades performance in diverse non-IID scenarios.
@@
Our experiments demonstrate that SCARLET consistently outperforms other distillation-based FL methods, achieving target accuracy with reduced communication and faster convergence. This contribution establishes SCARLET as a promising approach for efficient and scalable FL across heterogeneous client environments.
@@
  • Development of Enhanced ERA: We propose Enhanced ERA, an aggregation mechanism that replaces the mathematically unstable, temperature-based tuning of conventional ERA with an intuitive and provably robust, ratio-based control over aggregation sharpness, ensuring stable performance across diverse data distributions from strong non-IID to IID.
@@ II-A Federated Learning (FL)
Federated Learning (FL) is a framework for training machine learning models across decentralized clients while preserving data privacy. Since its introduction by [fedavg] with the Federated Averaging (FedAvg) algorithm, FL research has focused on challenges like communication efficiency, non-IID data distributions, and scalability.
Handling heterogeneous data, or non-IID data, where client distributions differ significantly, is a key challenge in FL. Traditional FL methods such as FedAvg are not optimized for these non-IID settings, resulting in slower convergence and reduced accuracy. Various solutions have been proposed, including personalized FL methods like pFedMe [pfedme], which allows client-specific adaptation, and regularization-based approaches such as FedProx [fedprox], which adds a penalty term to local updates to mitigate client drift. SCAFFOLD [scaffold] addresses local model divergence using control variates, and FedNTD [fedntd] employs selective distillation to retain knowledge from “not-true” classes, maintaining a global perspective without extra communication.
@@ II-B Communication-Efficient FL
Quantization reduces the bit precision of model updates, thus lowering data size. FedPAQ [fedpaq] applies quantization to compressed client updates for efficiency. FedCOMGATE [fedcomgate] combines quantization with gradient tracking, aligning local and global updates to support non-IID data while reducing communication.
Sparsification transmits only a subset of gradient elements, focusing on key information. Adaptive Gradient Sparsification [ags] dynamically adjusts sparsity based on client needs, using a bidirectional top-k approach that retains key gradients to balance communication efficiency and model accuracy.
Knowledge distillation, which shares distilled knowledge (such as output logits) instead of full parameters, allows model heterogeneity by transmitting only essential information. This approach accommodates model heterogeneity among clients by transferring essential output information rather than the model structure itself. Within distillation methods, approaches like FedDF [feddf] aggregate client outputs on the server, allowing clients to download an updated model without the need to exchange full parameters. Similarly, FedKD [fedkd] reduces communication costs by employing mutual distillation between a global model and a local model while dynamically compressing exchanged gradients, achieving significant communication savings without sacrificing performance. However, our research focuses on a different class of distillation methods that exclusively exchange model outputs bidirectionally between clients and the server, further enhancing communication efficiency.
@@ II-C Distillation-Based FL
Distillation-based FL reduces the communication overhead from parameter exchange by transmitting only model outputs (logits or soft-labels). Unlike traditional knowledge distillation [kd], which compresses models, distillation-based FL enables collaborative learning among heterogeneous clients, managing model heterogeneity while reducing communication.
Some studies use public data in distillation. FedMD [fedmd] enables clients with different models to collaborate by sharing logits derived from a labeled public dataset, allowing clients to refine models without sharing raw data. FD [fd] reduces communication by aggregating label-wise average soft-labels from clients’ private data, while DS-FL [dsfl] improves on FD by using unlabeled public data and introducing Entropy Reduction Aggregation (ERA) to handle high-entropy aggregated soft-labels in non-IID settings.
Recent methods further optimize communication, personalization, and non-IID handling. CFD [cfd] employs soft-label quantization and delta coding to reduce data transmission, and COMET [comet] clusters clients with similar data, sharing only soft-labels for improved personalization. Selective-FD [selectivefd] filters ambiguous knowledge through selective sharing, using client-side selectors and a server-side selector to handle out-of-distribution samples and high-entropy predictions.
However, existing distillation-based FL methods suffer from two fundamental limitations. First, they involve redundant communication, as they do not evaluate how significantly soft-labels are updated in each round. Second, their aggregation mechanisms (e.g., conventional ERA) are often mathematically unstable, leading to poor performance in non-IID settings. These limitations motivate our proposed method, SCARLET, which explicitly addresses such inefficiencies by introducing both a soft-label caching mechanism and the stable Enhanced ERA mechanism.
@@ II-D FL with Caching Mechanisms
Several studies have proposed FL methods incorporating caching mechanisms. CacheFL [cachefl] employs a caching mechanism that locally stores global models to reduce training time, enabling resource-limited clients to quickly access cached models. FedCache [fedcache] utilizes sample-specific hash mappings and server-side caching to efficiently deliver personalized knowledge without relying on public datasets. FedCache 2.0 [fedcache2] replaces logit exchanges with dataset distillation, treating the distilled datasets as a centralized knowledge cache on the server, from which clients download information aligned with their data.
@@ III-A Background and Notations
In FedAvg [fedavg], a standard FL method, each client $k\in[K]$ first updates its model $\mathbf{w}_{k}$ with its private dataset $\mathcal{B}_{k}$ using stochastic gradient descent (SGD). Then, each client uploads its updated model parameters or gradients to the server. The server aggregates these uploaded parameters to obtain the updated global model $\mathbf{w}_{g}$ as follows:
@@
where $\lvert\cdot\rvert$ denotes the cardinality (size) of a set. Subsequently, the server broadcasts the global model $\mathbf{w}_{g}$ to all clients. This procedure is repeated for a finite number of communication rounds.
In DS-FL [dsfl], a distillation-based FL method that serves as the baseline for SCARLET, each client and the server exchange soft-labels instead of model parameters. We assume that each client $k\in[K]$ has not only its labeled private dataset $\mathcal{B}_{k}$ but also the public unlabeled dataset $\mathcal{P}$. After each client updates its model $\mathbf{w}_{k}$ as in FedAvg, it generates local soft-labels $z_{k}^{t}=\sigma(\mathbf{x};\mathbf{w}_{k})$ for $\mathbf{x}\in\mathcal{P}^{t}$, where $\mathcal{P}^{t}$ is unlabeled public dataset used in communication round $t$. The local soft-labels $z_{k}^{t}$ are uploaded from each client to the server, and the server aggregates them to create the global soft-labels $\hat{z}^{t}$ with ERA as follows:
@@
where $\eta_{\mathrm{dist}}$ is the learning rate and $\phi_{\mathrm{dist}}(\cdot,\mathbf{w}_{k}^{t},\cdot)$ is the loss function in the distillation procedure. The loss function $\phi_{\mathrm{dist}}$ is exemplified by the Kullback–Leibler divergence. Along with the clients’ local models, the global model $\mathbf{w}_{g}^{t}$ on the server is also updated with the global soft-labels $\hat{z}^{t}$ and public unlabeled dataset $\mathcal{P}^{t}$.
To avoid ambiguity, we clarify that in this paper, we refer to the outputs shared in distillation-based FL methods as soft-labels rather than logits. Although the term logits is often used in some prior works, such as [dsfl], [fd], to refer to the output of the softmax function, it more commonly denotes the unnormalized model output before softmax function. Consequently, we will consistently use the term soft-labels to describe the normalized outputs used in the distillation process, aligning with standard terminology in classification tasks.
@@
NN Number of classes for the classification task
@@
Algorithm 1 SCARLET (Full Participation Scenario)
1:Server Initialization:
2:Initialize global model $\mathbf{w}_{g}$ and all client models $\{\mathbf{w}_{k}\}_{k=1}^{K}$
3:Initialize global cache $\mathcal{C}_{g}$ and all local caches $\{\mathcal{C}_{k}\}_{k=1}^{K}$
4:Set cache duration $D$
5:Distribute public dataset $\mathcal{P}$ to all clients $k\in[K]$
6:
7:for $t=1,\cdots,T$ communication rounds do:
8:  Server:
9:  Randomly select subset $\mathcal{P}^{t}\subset\mathcal{P}$
10:  Compute $\mathcal{I}_{\text{req}}^{t}\leftarrow\{i\in\mathcal{I}^{t}\mid\mathcal{C}_{g}(i)\text{ does not exist}\}$
   $\triangleright$ Identify samples not cached to request
11:  Send $\mathcal{I}_{\text{req}}^{t}$ to all clients $k\in[K]$
12:
13:  if $t>1$ then
14:   Send $\gamma^{t-1},\hat{z}_{\text{req}}^{t-1},\mathcal{I}^{t-1}$ to all clients $k\in[K]$
15:  end if
16:
17:  Client $k\in[K]$ (in parallel):
18:  $\mathbf{w}_{k}^{(t,0)}\leftarrow\mathbf{w}_{k}^{(t-1,E)}$
19:  if $t>1$ then
20:   $\hat{z}^{t-1}\leftarrow$
     UpdateLocalCache($\mathcal{C}_{k}$, $\gamma^{t-1}$, $\hat{z}_{\text{req}}^{t-1}$, $\mathcal{I}^{t-1}$)
      $\triangleright$ See Algorithm 2 for definition
21:   $\mathbf{\tilde{w}}_{k}^{(t,0)}\leftarrow\mathbf{w}_{k}^{(t,0)}$
22:   for $i=1,\cdots,E_{\text{dist}}$ do: $\triangleright$ Distillation step
23:     $\mathbf{\tilde{w}}_{k}^{(t,i)}\leftarrow\newline \hskip 60.00009pt\mathbf{\tilde{w}}_{k}^{(t,i-1)}-\eta_{\text{dist}}\nabla\phi_{\text{dist}}(\mathcal{P}^{t-1},\mathbf{\tilde{w}}_{k}^{(t,i-1)},\hat{z}^{t-1})$
24:   end for
25:   $\mathbf{w}_{k}^{(t,0)}\leftarrow\mathbf{\tilde{w}}_{k}^{(t,E_{\text{dist}})}$
26:  end if
27:  for $j=1,\cdots,E$ do: $\triangleright$ Local training step
28:   $\mathbf{w}_{k}^{(t,j)}\leftarrow\mathbf{w}_{k}^{(t,j-1)}-\eta\nabla\phi(\mathcal{B}_{k},\mathbf{w}_{k}^{(t,j-1)})$
29:  end for
30:  $\mathcal{P}_{k}^{t}\leftarrow\{\mathcal{P}[i]\mid i\in\mathcal{I}_{\text{req}}^{t}\}$
31:  Send soft-labels $z_{k}^{t}\leftarrow\sigma(\mathcal{P}_{k}^{t},\mathbf{w}_{k}^{(t,E)})$ to server
32:
33:  Server:
34:  Set $\mathbf{w}_{g}^{(t,0)}\leftarrow\mathbf{w}_{g}^{(t-1,E)}$
35:  Aggregate $\{z_{k}^{t}\}_{k\in[K]}$ with Enhanced ERA to get $\hat{z}_{\text{req}}^{t}$
       $\triangleright$ See (Eq. 4)
[…]
37:  $\hat{z}^{t}\leftarrow\hat{z}_{\text{req}}^{t}\cup\hat{z}_{\text{cached}}^{t}$ $\triangleright$ Assemble full soft-label set
38:  for $i=1,\cdots,E_{\text{dist}}$ do: $\triangleright$ Global distillation step
39:   $\mathbf{w}_{g}^{(t,i)}\leftarrow\mathbf{w}_{g}^{(t,i-1)}-\eta_{\text{dist}}\nabla\phi_{\text{dist}}(\mathcal{P}^{t},\mathbf{w}_{g}^{(t,i-1)},\hat{z}^{t})$
40:  end for
41:  $\gamma^{t}\leftarrow$ UpdateGlobalCache($\mathcal{C}_{g}$, $\hat{z}^{t}$, $\mathcal{I}^{t}$, $t$, $D$)
      $\triangleright$ See Algorithm 2 for definition
42:end for